OpenStudy (anonymous):

Solve the equation: 9x − 2 = 4(x + 3)

OpenStudy (anonymous):

Distribute the four: 9x - 2 = 4x +12 Subtract 4x from both sides 5x - 2 = 12 Add 2 to both sides: 5x = 14 Divide both sides by 5: x = 14/5

OpenStudy (saifoo.khan):

I agree with Hero.

OpenStudy (anonymous):

thanksss for helping

OpenStudy (anonymous):

You're most welcome